OpenStudy (anonymous):

erica bought 3 1/2 yards of fabric. if she uses 2/3 of the fabric to make a curtain, how much will she have left?

OpenStudy (anonymous):

1/3*? what

OpenStudy (anonymous):

\[3\frac{1}{2} = \frac{7}{2}\] \[\frac{7}{2} \times \frac{1}{3} = ?\]